Given this information for a monopoly company producing cars :


  1. cost of production : 2 million ( 2000 in thousands) w/ nothing to maintain 
  2. foreseeable demand : q= 900-150p ( q is the number of drivers (thousands) given the price per driver p)


What would be the profit-maximizing problem (optimization)


What would be the profit-maximizing quantity and price?


Thank you in advance (:


1
Expert's answer
2021-02-22T14:03:17-0500

The profit maximizing problem for monopolist is to produce the quantity which provides equality of marginal cost and marginal revenue. MR=MC. If it costs nothing to maintain, MC=0. MR=pq'=p(900-150p)'=900-300p. So, 900-300p=0, then p=3, q=900-150×3=450.